OpenStudy (cutiekawdfi):

help? solve by factoring x^2 = 4 + 12

OpenStudy (anonymous):

first simplify the right side of equation

OpenStudy (anonymous):

what u will get?

OpenStudy (anonymous):

the equation will look like this: x^2=16 now, since you need to solve by factoring, subtract 16 from both sides. x^2 - 16=16 - 16 x^2 - 16 = 0 @cutiekawdfi

OpenStudy (cutiekawdfi):

i messed up the equation.. its actually x^2= 4x + 12

OpenStudy (anonymous):

ok. then do the following subtarct 4x and 12 from both sides you will get x^2 - 4x - 12 = 0 now you need to get such equation: (x-a)*(x-b)=0, where x=a and x=b will be your solutions.

OpenStudy (anonymous):

so you need to find such a and b, so that a+b=-4 and a*b = -12

OpenStudy (anonymous):

oops sorry a+b=4 and a*b=-12

OpenStudy (anonymous):

so it will be a=6 and b=-2 6+(-2)=4.......... this one works 6*(-2)=-12............this one also works so (x-6)(x-(-2))=0 or (x-6)(x+2)=0 so the solutions are 6 and -2